OpenStudy (anonymous):

-2+(4-5)2-8/2 evaluate

OpenStudy (anonymous):

thats a square (2)

OpenStudy (anonymous):

You mean (4-5) the whole squared?

OpenStudy (anonymous):

i mess up hold on

OpenStudy (anonymous):

-2+3(4-5)2(sq)-8/2

OpenStudy (anonymous):

-2 + 3 - 8/2 = 1 - 8/2 = -6/2 = -3